OpenStudy (anonymous):

Which equation is not equivalent to the formula m = ca? a=m/c a=c/m m=ac c=m/a

OpenStudy (anonymous):

a = c/m If you rewrite this you end up with m=c/a

OpenStudy (anonymous):

b

OpenStudy (lilg132):

a = c/m

OpenStudy (lilg132):

no?

OpenStudy (domebotnos):

what is it?